Poisson Distribution (Cambridge (CIE) A Level Maths: Probability & Statistics 2): Exam Questions

Exam code: 9709

4 hours27 questions
1a
2 marks

A random variable X~Po(5).

Write down the integer values of a and b which make the following statements true.

(i) P(X≤7)=1−P(X≥a)

(ii) P(X≥3)=1−P(X≤b)

1b
6 marks

Find, giving your answers correct to 3 significant figures,

(i) P(X=4)

(ii) P(X≤2)

(iii) P(X≥2)

2
5 marks

A random variable X~Po(3).

Find the exact value, in terms of e, of

(i) P(X=5)

(ii) P(X≤1)

(iii) P(X≥2)

3a
3 marks

The random variable X~Po(λ).

Find the value of P(X=0) in the case when

(i) λ=7

(ii) λ=2.5

(iii) λ=0.01

3b
3 marks

Find the value of P(X=1) in the case when

(i) λ=7

(ii) λ=2.5

(iii) λ=0.01

3c
2 marks

Given that λ=k, find a simplified expression in terms of k for

(i) P(X=0)

(ii) P(X=1)

4
3 marks

For each of the following random variables, find the mean and the standard deviation.

(i) X~Po(4)

(ii) X~Po(1)

(iii) X~Po(5.4)

5a
3 marks

The random variable X~Po(6.3).

(i) By taking out a common factor of e−λ, factorise the expression

e−λλ44!+e−λλ55!+e−λλ66!+e−λλ77!+e−λλ88!

(ii) Hence calculate P(4≤X≤8).

5b
4 marks

Calculate

(i) P(X≤3)

(ii) P(5≤X≤7)

6a
4 marks

A random variable X~Po(6).

Find the exact value, in terms of e, of

(i) P(X=3)

(ii) P(X=4)

(iii) P(X=5)

6b
2 marks

Use your answers to part (a) to find the exact value of

(i) 64P(X=3)

(ii) 65P(X=4)

6c
1 mark

Compare your answers to part (a) and part (b), then use the pattern to write down an expression, in terms of k, for

6k+1P(X=k)

where k is an integer and k≥0.

7a
2 marks

Naomi, a teacher, receives emails at a constant average rate during the school day. Naomi models the number of emails she receives within a fixed period of time using a Poisson distribution, Po(λ).

Write down two assumptions that Naomi has made about the emails in order to use a Poisson model.

7b
3 marks

Naomi knows that on average she receives 10 emails in one hour.

Calculate the value of λ in the case when Naomi is modelling the number of emails she receives in

(i) two hours

(ii) 15 minutes

(iii) one hour and 10 minutes.

8a
2 marks

The random variables R~Po(2), S~Po(3) and T~Po(4) are independent.

Write down the distribution of

(i) R+S

(ii) S+T

8b
3 marks

(i) Write down the distribution of R+T.

(ii) Find P(R+T=5).

9a
2 marks

The number of times that the internet connection at Amira's house disconnects during a day is denoted by X. On average, the internet disconnects 5 times each day.

State a suitable distribution for X, giving the value of any parameter, and state one assumption that is required for this model to be valid in this context.

9b
6 marks

Find the probability that, during a randomly chosen day, the internet at Amira's house disconnects

(i) exactly 4 times

(ii) at most 3 times

(iii) at least 2 times.

10a
1 mark

The numbers of pieces of ham and pieces of salami on a pizza from Pizza Palace have the independent distributions Po(6.2) and Po(4.9) respectively. Roberto orders a pizza from Pizza Palace.

State the distribution of the total number of pieces of meat on Roberto's pizza, giving the value of any parameter.

10b
2 marks

Find the probability that Roberto's pizza contains a total of exactly 10 pieces of meat.

10c
2 marks

Find the probability that Roberto's pizza contains more than 9 and fewer than 13 pieces of meat.

1a
2 marks

The number of times that Lucy goes to the cinema in a week has a Poisson distribution with mean 4.

Find the probability that Lucy goes to the cinema exactly 5 times in a randomly chosen week.

1b
3 marks

Find the probability that Lucy goes to the cinema not more than 4 times in a randomly chosen 2-week period.

1c
3 marks

Find the probability that Lucy goes to the cinema at least once on a randomly chosen day.

2a
3 marks

Customers enter Comic Stans, a comic book store, at random and independently at a constant average rate of 8 every 15 minutes.

Find the probability that exactly 3 customers enter the store during a randomly chosen 1-minute period.

2b
3 marks

Find the probability that at least 1 customer enters the store during a randomly chosen 15-second period.

2c
3 marks

Find the probability that at most 3 customers enter the store during a randomly chosen 10-minute period.

2d
1 mark

Find the variance of the number of customers entering the store during a 1-hour period.

3a
2 marks

Amber suggests that the number of times that she hiccups can be modelled by a Poisson distribution.

State two conditions that must be satisfied for a Poisson distribution to be a suitable model in this context.

3b
8 marks

The mean number of hiccups in a 30-second period is 2.9. Assume now that a Poisson distribution is a suitable model.

Find the probability that Amber hiccups

(i) exactly 3 times in a randomly chosen 30-second period

(ii) at least twice but not more than 5 times in a randomly chosen 15-second period

(iii) at least once in a randomly chosen 1-minute period.

4a
5 marks

The table below shows the data from a sample of 50 observations of a variable x.

(i) Calculate unbiased estimates for the mean and the variance.

(ii) State, with a reason, whether a Poisson distribution could be used to model the population’s data.

x

0

1

2

3

4

Frequency

3

5

17

15

10

4b
5 marks

The table below shows the data from a sample of 100 observations of a variable y.

(i) Calculate unbiased estimates for the mean and the variance.

(ii) State, with a reason, whether a Poisson distribution could be used to model the population’s data.

y

0

1

2

3

4

5

Frequency

15

19

25

21

12

8

5a
4 marks

Jim records the number of birds that fly past his window during each of 120 randomly chosen 10-minute periods. The results are summarised in the table.

Number of birds

Frequency

0

43

1

44

2

22

3

8

4

3

5 or more

0

Calculate unbiased estimates of the population mean and variance of the number of birds that fly past Jim's window in a 10-minute period.

5b
1 mark

Explain how your answers to part (a) suggest that a Poisson distribution may be a suitable model for the number of birds that fly past Jim's window in a 10-minute period.

5c
9 marks

Jim models the number of birds that fly past his window in a 10-minute period by the distribution Po(1).

Use Jim's model to find the probability that

(i) exactly 2 birds fly past Jim's window in a 30-minute period

(ii) fewer than 2 birds fly past Jim's window in a 1-minute period

(iii) at least 4 birds fly past Jim's window in a 1-hour period.

6a
2 marks

The number of times that Blue, a dog, barks is modelled by a Poisson distribution with mean 25 per hour.

State two assumptions that are required for the Poisson model to be valid in this context.

6b
9 marks

Find the probability that Blue barks

(i) exactly 5 times in a randomly chosen 10-minute period

(ii) at most 4 times in a randomly chosen 15-minute period

(iii) more than 47 times but not more than 51 times in a randomly chosen 2-hour period.

7
6 marks

Roger is investigating historical data from World War II. In south London, an area of 144 km² was divided into 576 equal squares, and the number of times each square was hit by a flying bomb over a period of time was recorded. During this time, a total of 537 bombs landed in the area. The results are shown in the table below.

Number of times hit by a flying bomb

Number of squares

0

229

1

211

2

93

3

35

4

7

5 or more

1

Roger suggests that a Poisson distribution with mean 537576 would be an appropriate model for the number of times a square was hit by a flying bomb.

(i) By finding probabilities using Roger's model, estimate the expected number of squares that would be hit, respectively, by 0, 1, 2, 3, 4 and 5 or more flying bombs.

(ii) By comparing the actual number of squares in the table with the expected numbers of squares from part (i), state whether Roger's model is appropriate.

1a
1 mark

The number of customer enquiries received per hour at an office is denoted by X. It is given that X has a Poisson distribution with standard deviation 2.25.

Write down the value of E(X).

1b
2 marks

Find P(X=4).

1c
3 marks

Find P(1.5≤X<5).

1d
3 marks

Find the probability that X=4 given that X>0.

2a
3 marks

Grace attends nursery 5 days a week. The number of tantrums that Grace has in a day at nursery is modelled by a Poisson distribution with variance 3.14.

Find the probability that Grace has exactly 17 tantrums during a randomly chosen week at nursery.

2b
3 marks

Find the probability that Grace has fewer than 4 tantrums in a randomly chosen 2-day period at nursery.

2c
4 marks

Given that Grace has fewer than 4 tantrums at nursery on a randomly chosen day, find the probability that she has no tantrums at nursery on that day.

3a
5 marks

The numbers of detentions issued per school day by Ms Ottway, a teacher, over a period of 150 days are summarised in the table.

Number of detentions

Number of days

0

51

1

54

2

36

3

6

4

3

5 or more

0

Calculate unbiased estimates of the population mean and variance of the number of detentions issued per day, and explain how these estimates suggest that a Poisson distribution is an appropriate model.

3b
2 marks

Using a Poisson distribution with mean equal to your estimate from part (a), find the probability that Ms Ottway issues at least 5 detentions on a randomly chosen day.

3c
4 marks

Students at Ms Ottway's school attend school 5 days a week for 40 weeks in a year.

Estimate the number of weeks in a school year in which Ms Ottway issues fewer than 3 detentions.

4
7 marks

X, Y and Z are random variables with X~Po(α), Y~Po(β) and Z~Po(γ).

(i) Given that P(X=3)=P(X=7), find the value of α.

(ii) Given that P(Y=0)=0.253, find the value of β.

(iii) Given that 5Var(Z)=(E(Z))2−6, find the value of γ.

5a
3 marks

In a computer game, enemies appear at random and at a constant average rate of 2.5 per minute. You may assume that enemies appear independently of one another.

Find the probability that exactly 10 enemies appear during a randomly chosen 5-minute period.

5b
3 marks

Find the probability that at least 3 enemies appear during a randomly chosen 90-second period.

5c
3 marks

The probability that at least one enemy appears during a k-minute period is 0.999.

Find the value of k.

6a
4 marks

Matt receives messages from three friends: Danny, Jessica and Luke. On average, Matt receives 1.7 messages per hour from Danny, 1.5 messages per hour from Jessica and 1.1 messages per hour from Luke.

Stating a necessary assumption, find the probability that Matt receives fewer than 4 messages from Danny during a randomly chosen 2-hour period.

6b
3 marks

Stating a further assumption needed, find the probability that Matt receives exactly 6 messages in total during a randomly chosen 1-hour period.

6c
2 marks

Find the probability that Matt receives at least one message from any of the three friends during a randomly chosen 1-hour period.

6d
3 marks

Find the probability that Matt receives at least one message from each friend during a randomly chosen 1-hour period.

7a
2 marks

Phoebe has a faulty smoke detector which rings at an average rate of 5 times per day.

State two conditions necessary for a Poisson distribution to be a suitable model for the number of times that the smoke detector rings.

7b
6 marks

Assume now that a Poisson distribution is a suitable model.

(i) Find the probability that the smoke detector rings exactly 3 times on each day of a randomly chosen 4-day period.

(ii) Find the probability that the smoke detector rings exactly 12 times during a randomly chosen 4-day period.

(iii) Explain why the answers to part (b)(i) and part (b)(ii) are different.

7c
4 marks

Given that the smoke detector rings at least once during a randomly chosen 6-hour period, find the probability that it rings no more than 4 times during this period.

8
7 marks

The random variables X~Po(α), Y~Po(β) and Z~Po(γ).

(i) Given that P(X=4)=P(X=2)+P(X=3), find the value of α.

(ii) Given that P(Y≥1)=0.314, find the value of β.

(iii) Given that P(Z=k+1)=P(Z=k−1) where k is a positive integer, find an expression for γ in terms of k.

1a
4 marks

The number of mistakes made by a student, Priya, in a 20-minute revision period is modelled by the distribution Po(1.2). The number of mistakes made by a different student, Qays, in a 30-minute revision period is modelled by the distribution Po(2.2).

Stating a necessary assumption, find the probability that, in a randomly chosen 1-hour revision period, Priya makes exactly 2 mistakes and Qays makes exactly 1 mistake. Give your answer in the form abec, where a, b and c are integers.

1b
4 marks

Given that Priya and Qays make a total of exactly 9 mistakes in a randomly chosen 1-hour revision period, find the probability that Priya makes exactly 5 mistakes.

1c
2 marks

Given instead that Priya makes exactly 5 mistakes in a randomly chosen 1-hour revision period, find the probability that Priya and Qays make a total of exactly 9 mistakes in that period.

1d
3 marks

In the week before a test, Priya has ten 30-minute revision periods.

Estimate the number of these revision periods during which Priya makes at least one mistake.

2a
4 marks

Spelling mistakes in essays written by Gamu occur at random and at a constant average rate of 7 per 150 words. Gamu models the number of spelling mistakes using a Poisson distribution.

Find the largest number of words that Gamu can write so that the probability that she makes at least one spelling mistake does not exceed 0.75.

2b
6 marks

Gamu writes three short essays, containing 100 words, 200 words and 250 words. An essay is returned by her lecturer if more than 1% of its words contain spelling mistakes.

You should assume that each spelling mistake is in a different word, and that the numbers of spelling mistakes in the three essays are independent.

Find the probability that all three essays are returned.